Shift Summary Day 3 Aug., 1999

Christine Pietrykowski (pietryko@hep.uchicago.edu)
Tue, 3 Aug 1999 15:28:44 -0500

Shift Summary for Day shift, 08/3/99

On Shift: Nick, Christine Pietrykowski

Hours of beam available at all intensities: 7.75
Hours of beam available at slightly below request intensity: 7.75
Hours of E832 data (special runs not included): 7.25
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Changing tapes, and a Drt Fresh start., 45
min where there was no beam.

========================================================================
Runs ended this shift: 14092 this was still the special lower
intensity with 2.e12 protons on target and unlimited ke3s,
and 14093 ktev3 have a cpu problem, needed drt fresh start.
========================================================================
Tapes brought to FCC this shift: QKE010-018, QKB073
========================================================================
Accesses during shift: None
========================================================================
Overview of shift: Data taking, changing tapes.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14092 377 7:00 New shift takes over, No beam.
378 7:01 Beam Returns- 1.99e12
485 9:22 Tapes full, run ends.
9:30 Changed the analysis magnet to positive.
14093 9:53 New run starts. Back to normal tapes.
Going back to nomininal beam, 4.2e12, and
normal e832_beam triggers.

23 Beam went away-TeV RF station 8 trips, est. 5
min.
34 10:37 Beam returns.
52 Run ended due to ktev3 CPU dead.
14094 11:15 Run starts
9 11:30 Called MCR to check on getting higher beam
intensity, it was a success! Intensity went
from 3.8 to 4.01
126 14:05 KTeVana errors, ran kfrend.
140 14:24 Pipeline errors, reset pipeline
141 14:26 KTeVana errors, ran kfrend.
146 KTeVana errors, ran kfrend.
150 14:36 Beam went away- DO low beta QPM problems, est.
10 min.
153 14:41 Beam returns, read out errors.
Pipeline errors, reset pipeline.
Ran kfrend.
155 Good spill with beam.
163 14:54 Beam went away- Rebooting low beta QPM, est
10 min.
15:30 Still no beam, awaiting its return.

========================================================================
Follow up Notes on Previous Shift: Keep taking data if possible.
========================================================================
Items to Follow up on Next Shift: Keep taking data if possible.